### v1

# Relations Among Different Inequality Measures in Complex Systems: From Kinetic Exchange to Earthquake Models

Abstract
We present a numerical study of several inequality measures across two kinetic wealth-exchange models with extreme inequality features (namely the Banerjee model, and the Chakraborti or Yard-Sale model) and two earthquake simulating models (namely the Chakrabarti-Stinchcombe two-fractal overlap model and the nonlinear dynamical Burridge-Knopoff model), and a synthetic Pareto distribution. For each model we compute numerically the Lorenz function for the respective models’ wealth, overlap magnitude or avalanche distributions. We then estimate the variations of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices in these models with systematic variations of saving propensity (for the two wealth-exchange models), with systematic variations of generation or block numbers (for the two earthquake simulating models). We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models and the two-fractal overlap model) remain a little above unity (theoretically predicted value) and deviating a little higher by a maximum of $4\%$ near $g=k\simeq 0.86$, which was identified earlier to be the precursor point of criticality in several self-organized critical models ($k=0.80$ corresponds to Pareto’s 80-20 law). In the Burridge-Knopoff model for some instances of time, the value of p/(2k1)p/(2k-1) drops a little below unity. This and some other quantitatively similar behaviors of the inequality indices across socio-economic and geophysical models may provide a coherent and comparative framework for identifying the subtle features in the statistics of such disparate dynamical systems.
keywords:
Inequality , Kinetic wealth exchange models , Pareto distribution , Earthquake models , Gini index , Pietra index , Kolkata index
[…]
1 Introduction
Inequality arises not only in socio-economic systems but also in a wide range of physical and geophysical processes. To compare these behaviors within a unified framework, we numerically quantify inequality in four classes of models: two kinetic wealth-exchange models, a synthetic Pareto distribution, and two earthquake-simulating models.
For the socio-economic domain, early models introduced by Dragulescu and Yakovenko [DY_2000] and Chakraborti and Chakrabarti [CC_2000] demonstrated that simple stochastic exchange rules can generate realistic wealth distributions. Extensions involving fixed or distributed saving propensity [CCM_2004], interaction heterogeneity, and asymmetric exchange have yielded richer phenomenology. The Chakraborti Model [C_Model] or also known as Yard-Sale model [YS], later extended by Boghosian and collaborators [YS1], captures multiplicative dynamics and extreme wealth condensation. The Banerjee Model [B_Model, B_Model1] represents another natural kinetic exchange framework, whose intrinsic trading dynamics lead to steady-state configurations characterized by extreme concentration of wealth among a small subset of agents. In this study, we focus on the Banerjee Model, where inequality is tuned by varying the range of exchange and the saving propensity, and the Chakraborti Model (or Yard-Sale model), where saving propensity acts as the central control parameter.
To broaden the scope beyond socio-economic systems, we also examine inequality associated with Pareto-type distributions [Pareto, Pareto1] by tuning their shape parameter $\alpha$, given the ubiquity of power laws in systems exhibiting self-organized criticality. Since such distributions are central to geophysical phenomena, we analyze inequality patterns in two different earthquake models, the Chakrabarti–Stinchcombe two-fractal overlap model [CS_Model, CS1], where event sizes arise from Cantor-set overlaps, and the nonlinear Burridge–Knopoff spring-block model [langer_1994, BK_Model], where block–spring dynamics generate heterogeneous slip events.
Quantifying inequality across these systems , we therefore investigate three Lorenz curve [lor, lor1] based measures: the Gini index (gg[gini], captures the average deviation from perfect equality; The Pietra (or Hoover) index (pp[pietra, pietra1], measures the maximum imbalance in cumulative wealth and the more recently introduced Kolkata index (kk[k-index, k-index1, k-index2, k-index3], identifies the wealth share of the richest fraction of the population and has found applications in diverse socio-economic contexts. For each system, we systematically vary the relevant control parameters such as the exchange range ($R$) and saving propensity ($\lambda$) in the Banerjee Model, saving propensity ($\lambda$) in the Chakraborti Model, the exponent ($\alpha$) in the synthetic Pareto distribution, the Cantor generation index in Chakrabarti–Stinchcombe two-fractal overlap model, and the time evolution of the Burridge–Knopoff model and compute the corresponding inequality indices. In a recent work [asim_2025] authors studied the ratios p/(2k1)p/(2k-1) and $p/g$ in some financial systems. We further examine the ratios p/(2k1)p/(2k-1), $p/g$ for these kinetic exchange models, the two earthquake simulating models, and the Pareto distribution to investigate the variation of these and other relationships among the inequality measures $g$, $p$ and $k$.
Together, these analyses provide a unified quantitative perspective on how inequality evolves across these models that governed by very different microscopic rules. We can see how tuning a small number of parameters systematically shapes these outcomes. The paper is organized as follows: Section 2 introduces the Lorenz curves and the inequality measures investigated in this study. Section 3 outlines the four models and the corresponding simulation framework. Section 5 presents the numerical results and analysis. Finally, Section 6 offers a brief summery of the main results obtained and the concluding remarks.
2 Inequality Measures
Quantifying inequality in any distribution whether describing wealth, event sizes, or released energy (hereafter called wealth) requires statistical measures that capture heterogeneity in a systematic manner. In this study, we numerically examine three statistical measures based on the Lorenz function [lor, lor1], namely the Gini index [gini], the Pietra (or Hoover) index [pietra, pietra1], and the Kolkata index [k-index, k-index2].
The Lorenz function (or Lorenz curve) $L(x)$ is a foundational tool for visualizing and quantifying inequality. From this function many inequality measures are derived. For a non–negative wealth $y$, if its distribution be $P(y)$, the Lorenz function is defined as:
$L(x)=\frac{\displaystyle\int_{0}^{x}y\,P(y)\,dy}{\displaystyle\int_{0}^{\infty}y\,P(y)\,dy},\qquad 0\leq x\leq 1.$ (1)
Here, $x$ denotes the bottom $x$ fraction of the population, and $L(x)$ gives the corresponding fraction of the wealth they hold. For a perfectly equal distribution, $L(x)=x$, producing a 45-degree line known as the perfect equality line (see in Fig. 1). Deviations of $L(x)$ below this line indicate inequality. If the total wealth goes to a single hand, we get the absolute inequality line as seen in Fig. 1.
2.1 Gini Index
The Gini index $g$ is one of the most widely used scalar measures of inequality, derived directly from the Lorenz curve. It ranges from 0 to $1$. From Fig. 1, one can see that $g$ measures the ratio of the area between the equality (blue) line and the Lorenz curve L(x) (orange curve) and the area between the equality (blue) line and the absolute inequality (red) line. In terms of the Lorenz curve:
$g=\frac{\int_{0}^{1}[x-L(x)]\,dx}{1/2}=1-2\int_{0}^{1}L(x)\,dx.$ (2)
(k,L(k))(k,L(k))kk(k,k)(k,k)(0,1)(0,1)(0,0)(0,0)(1,1)(1,1)(1,0)(1,0)ppL(x)L(x),  $L_{c}(x)$xxperfect equality lineLorenz curve L(x)L(x)absolute inequality linecomplementary Lorenz curve $L_{c}(x)$
Figure 1: Schematic diagram of Lorenz curve and inequality measures where x-axis represents cumulative fraction of population from poor to rich and y-axis depicts the cumulative fraction of their wealth. Gini index $g$ is given by the ratio of the area of yellow region and yellow++pink region. Pietra index $p$ is the maximum distance between the equality line and Lorenz curve and Kolkata index is the ordinate point $k$ where Lorenz curve cuts the off diagonal line (perpendicular to the equality line) at the point $(k,L(k))$.
2.2 Pietra Index
The Pietra index $p$ (also called the Hoover index or Robin Hood index) measures the maximum vertical distance between the Lorenz curve and perfect equality line (see Fig. 1). Equivalently, it represents the excess fraction of the wealth that must be redistributed from the richer people to the poorer people to achieve perfect equality. The Pietra index takes values in $[0,1]$, with higher values implying stronger inequality. It is defined as:
$p=\max_{0\leq x\leq 1}\left[\,x-L(x)\,\right].$ (3)
2.3 Kolkata Index
The Kolkata index (or $k$-index) is a more recent inequality measure, widely used in socio-economic systems and increasingly in complex physical systems. It is defined as the fixed point of the complementary Lorenz function ($L_{c}(x)$):
$L_{c}(k)\equiv 1-L(k)=k.$ (4)
In Fig. 1, from simple geometry one can see that $k$ fraction of the total wealth possessed by the richest (1k)(1-k) fraction of the population. Note that, $k=0.80$ implies that $20\%$ of the population holds $80\%$ of the wealth, reminiscent of the Pareto 80802020 rule. The $k$-index varies from $1/2$ (perfect equality) to $1$ (perfect inequality).
In many empirical systems and theoretical models, $k$ displays near-universal scaling behaviour in relation to $g$ or $p$, which motivates our numerically comparative study across two kinetic wealth exchange models and two earthquake simulating models [sand, manna, soumya, asim_2025].
3 Models
In this section, we describe four considered models for investigating inequality across these socio-economic and geophysical systems. Our analysis spans two kinetic wealth-exchange models, a synthetic Pareto distribution, and two earthquake-simulating models based on fractal overlap and spring–block dynamics.
3.1 Kinetic Wealth-Exchange Models
Kinetic wealth exchange models given by Dragulescu and Yakovenko [DY_2000] describes wealth evolution through pairwise stochastic exchanges among agents and Chakraborti and Chakrabarti through saving propensities [CC_2000]. Below we describes our considered two kinetic wealth exchange models.
3.1.1 B-Model with Exchange Range
The Banerjee model [B_Model, B_Model1] (or B-Model in short hereafter) introduces a constrained exchange rule in which each agent interacts only within a selected exchange range $R$ in the wealth-ordered list. At each time step $t$:
(i) Arrange agents by their current wealth from poor to rich.
(ii) Choose an agent $i$ at random.
(iii) Select the interaction or exchange partner $j$ randomly among the agents within range $R$ of $i$ in the wealth-ordered list.
The update rule reads:
$w_{i}(t+1)=\epsilon(w_{i}(t)+w_{j}(t)),\qquad w_{j}(t+1)=(1-\epsilon)(w_{i}(t)+w_{j}(t)),$ (5)
where $\epsilon$ is a random number drawn uniformly from $(0,1)$. This model naturally generates strong wealth condensation for small $R$, producing highly unequal steady states—though not to the extreme limit observed in the C-Model (discussed later), where all wealth collapses into a single agent. For each choice of $R$, we compute the steady-state Lorenz geometry and associated inequality indices after ensemble averaging.
3.1.2 B-Model with Savings
We also study a modified version of the Banerjee model in which agents retain a fixed fraction of their wealth before participating in the exchange. Each agent has a saving propensity $\lambda\in[0,1)$, so that only the remaining $(1-\lambda)$ portion of wealth is available for trading. The update rule becomes:
$w_{i}(t+1)=\lambda w_{i}(t)+\epsilon(1-\lambda)(w_{i}(t)+w_{j}(t)),$
$w_{j}(t+1)=\lambda w_{j}(t)+(1-\epsilon)(1-\lambda)(w_{i}(t)+w_{j}(t)),$ (6)
where $\epsilon\in(0,1)$ is a random sharing parameter. The introduction of saving stabilizes the dynamics and suppresses extreme condensation, producing a spectrum of inequality levels as $\lambda$ is varied. For each value of $\lambda$, ensemble-averaged steady-state distributions are used to compute the inequality measures.
3.1.3 C-Model with Savings
The Chakraborti Model [C_Model] (or C-Model in short hereafter) also widely known as Yard-Sale model [YS] is based on multiplicative dynamics in which the amount at stake depends on the wealth of the poorer agent. With saving propensity $\lambda\in[0,1)$, the effective tradable wealth becomes $(1-\lambda)\min(w_{i},w_{j})$. The exchanged amount is $\Delta w=\epsilon(1-\lambda)\min\{w_{i}(t),w_{j}(t)\}$, where $\epsilon$ is uniformly drawn from $(0,1)$. The update equations follow:
$w_{i}(t+1)=w_{i}(t)+\Delta w,\qquad w_{j}(t+1)=w_{j}(t)-\Delta w,$ (7)
with the sign depending on the random outcome of the exchange. This model is known for producing rich-get-richer dynamics and, for low saving propensities, may lead to complete condensation where all wealth eventually accumulates with a single agent. Ensemble-averaged steady-state distributions are used for computing all inequality metrics.
We have discussed about two kinetic wealth-exchange models, both of which generate exponential or gamma-like distributions. We now turn to earthquake simulation models, which yield power-law distributions. Although our focus here is not on analyzing the distributions produced by these models, but rather to investigate inequality within them. Before jumping from kinetic wealth exchange models to earthquake simulating models, we examine a synthetic Pareto distribution that exhibits different power-law behaviors as the exponent parameter ($\alpha$) is varied.
3.2 P-Model
We investigate inequality in synthetic Pareto-type distribution [Pareto, Pareto1] (or P-Model in short hereafter) characterized by the probability density function,
$p(x)=\frac{\alpha x_{m}^{\alpha}}{x^{\alpha+1}},\qquad x\geq x_{m},$ (8)
where $x_{m}$ is the minimum wealth (set to unity in our simulations) and $\alpha>0$ is the shape parameter. The parameter $\alpha$ controls tail heaviness: smaller $\alpha$ produces broader, more unequal distributions. The Lorenz curve takes the form,
$L(p)=1-(1-p)^{\frac{\alpha-1}{\alpha}},$ (9)
where $p$ denotes the cumulative fraction of the population, and $L(p)$ is the corresponding fraction of total wealth held by that population.
For each $\alpha$ we generate large synthetic datasets, compute the Lorenz curve, and evaluate the associated inequality indices. This baseline serves as a reference for comparing model-generated distributions with known power-law behavior observed in socio-economic and geophysical systems.
3.3 Earthquake Simulating Models
We now move to two classes of geophysical models that simulate earthquake like event size statistics. The Chakrabarti-Stinchcombe fractal overlap construction generates multifractal, heavy-tailed spatial measures that produce inequality through time variation of the overlap magnitude, whereas the Burridge–Knopoff model gives rise to inequality in terms of spatially varying spring force. Both models probe different but complementary aspects of earthquake dynamics, namely fault geometry and stress distribution during an earthquake. The Chakrabarti–Stinchcombe model exhibits a power-law scaling in the cumulative distribution of overlap magnitudes in the asymptotic limit of large generation number and large overlaps, arising purely from the construction of the underlying fractlas [pbhattacharya]. In contrast, the Burridge-Knopoff model displays power-law event-size statistics in specific dynamical regimes, where nonlinear friction and collective interactions drive the system toward scale-free behavior. Thus, although both models can produce similar phenomenological power-law statistics, the underlying mechanisms are fundamentally different– geometric hierarchy in the former and dynamical evolution in the latter.
3.3.1 CS-Model
The Chakrabarti Stinchcombe model [CS_Model, CS1, pbhattacharya] (CS-Model in short hereafter) based on the overlap of fractal fault surfaces (modeled using Cantor sets), focuses on fault geometry and the distribution of overlaps that correspond to stored elastic energy. It is therefore a model of geometric heterogeneity: it explains why slip events follow broad distributions (e.g., power laws) purely from the self-similar structure of the fault interface suitable for Lorenz-curve–based inequality analysis. Earthquake magnitudes empirically follow the Gutenberg–Richter law, $\log_{10}N_{r}(M>M_{0})=a-bM$, with $N_{r}(M>M_{0})$ denoting the number of events above magnitude $M_{0}$, and $b\simeq 1$. The model captures this scale-free behavior by treating a fault as two identical self-affine fractals, typically $n^{th}$-generation Cantor sets with $2^{n}$ segments of length $3^{-n}$, sliding over each other. A replica is shifted in discrete steps of $3^{-n}$ under periodic boundary conditions, and the overlap magnitude $Y_{n}(t)$ counts coinciding segments at step $t$, generating a discrete time series $\{Y_{n}(t)\}$. The allowed values of $Y_{n}$ are $2^{\,n-k}$ ($k=0,1,\dots,n$), occurring with exact frequencies
$N_{r}\!\left(2^{\,n-k}\right)=2^{k}\binom{n}{k}$ (10)
which normalize to the binomial probabilities $\Pr(2^{\,n-k})=\binom{n}{k}\left(\tfrac{1}{3}\right)^{n-k}\left(\tfrac{2}{3}\right)^{k}$; the most probable overlap scales as $\hat{Y}_{n}\sim(2^{n})^{1/3}$. CS-Model captures the static, structural statistics of fault roughness providing insight into how geometric irregularity shape the statistics of earthquakes.
3.3.2 BK-Model
In contrast, the Burridge–Knopoff Model [langer_1994, BK_Model] (BK-Model in short hereafter) is a dynamical and mechanical model of earthquake generation, where blocks connected by springs and subject to friction evolve over time. It is used to study stick–slip dynamics and stress redistribution. BK-Model captures the temporal dynamics of rupture propagation and seismicity. BK-Model provide insight into how dynamical frictional processes shape the statistics of earthquakes. This framework captures fault dynamics through a chain of $N$ blocks of mass $m$, each connected to its neighbors by springs of stiffness $k_{c}$ and to a slowly driven plate by springs of stiffness $k_{p}$. As the plate moves with velocity $v$, stress accumulates until blocks slip, producing earthquake-like events. Each block follows Newton’s equation of motion,
$m\ddot{u}_{i}=k_{c}\left(u_{i+1}-2u_{i}+u_{i-1}\right)+k_{p}\big(vt-u_{i}\big)-F_{\text{fric}}(\dot{u}_{i}),$ (11)
where the nonlinear friction force $F_{\text{fric}}(\dot{u_{i}})$ generates the characteristic stick–slip (SS) behavior, and is given by the expression:
$F_{\text{fric}}(\dot{u}_{i})$ $=\begin{cases}F_{s},&\text{if }\dot{u}_{i}=0\text{ and }|F_{\text{elastic}}|<F_{s},\\[5.0pt] \frac{1-\sigma}{1+\dfrac{2\mu\,v}{1-\sigma}},&\text{if sliding occurs},\end{cases}$ (12)
where $F_{s}$ is the static friction force $\mu$ is related to the friction weakening rate, and $\sigma$ is related to the friction drop parameter, when it changes from static to kinetic.
Using the stick–slip friction law—with parameters $\alpha$ (friction weakening rate) and $\sigma$ (friction drop), the model can reproduce Gutenberg–Richter–like magnitude statistics, exhibit chaotic dynamics despite being deterministic, and transition between periodic, chaotic, and scale-free regimes depending on parameters. Although more sophisticated rate-and-state (R&S) friction laws exist, we use the stick–slip rule which is computationally simpler.
4 Simulation Framework
For the both kinetic wealth exchange models: B-Model and C-Model describe above, we simulate the models with $N=100$ agents and perform $1000$ independent realizations for each parameter configuration using Eqns. (5), (3.1.2) and (7) respectively. Considering $w_{i}(t)$ be the wealth of agent $i$ at time $t$, the total wealth $W=\sum_{i=1}^{N}w_{i}(t)$ is always conserved. We initialize wealth as $w_{i}(0)=1$ for all $i$, and evolve the system for sufficiently long time ($t=10^{6}$) to ensure that a stationary distribution is reached. Inequality measures are computed from the steady-state wealth distributions averaged across realizations.
In the P-Model, agent incomes are drawn from a Pareto distribution defined on $1,\infty)$, with the minimum income set to unity and shape parameter $\alpha$, using Eqn. (9). Simulations are carried out for a population of $10^{6}$ agents and repeated over five independent realizations, with all reported quantities obtained by averaging over these realizations to reduce finite-size fluctuations.
For the CS-Model, the fractal representation of a fault surface is constructed using the classical base-3 Cantor set to model the geometrical overlap dynamics between two self-similar rough interfaces. For a given generation $n$, the frequency of each overlap magnitude is evaluated using Eqn. (10), with $n$ ranging from 4 to 200. The inequality measures $g$, $k$, and $p$ are computed from the corresponding Lorenz curves, where the overlap magnitude is plotted along the abscissa and the cumulative frequency along the ordinate. These measures provide a compact quantitative characterization of the degree of heterogeneity in the spatial overlap and enable us to track the evolution of inequality as the fractal resolution is progressively increased.
For the BK model, we perform numerical simulations to investigate the emergence of collective slip dynamics for a given set of physical parameters. The simulations are carried out under prescribed frictional and elastic conditions, specified by the friction coefficient $\mu$, the spring constants $k_{p}$, and the number of blocks $N$. At each time step, the instantaneous force acting on each block is computed using Eqns. (11) and (12), and the corresponding Lorenz curve is constructed by plotting the force magnitude along the x-axis, and the cumulative fraction of blocks along the y-axis. The degree of heterogeneity in the force distribution is quantified using inequality indices, which serve as compact descriptors of the non-uniformity in the stress landscape and provide a time-resolved measure of the buildup and release of inequality within the block ensemble.
5 Results And Numerical Analysis
5.1 B-Model with Exchange Range
In this study, our primary objective is to examine three well-known measures of economic inequality and to investigate how they relate to one another within the framework of the B-Model as we vary the exchange range parameter $R$. Using Eqns. (1)–(4), all three inequality indicators the Gini index ($g$), the Pietra index ($p$), and the Kolkata index ($k$) are derived from the Lorenz curve ($L(x)$), which characterizes the cumulative distribution of wealth across the population following Eqn. (5).
For each value of the exchange range parameter ($R$) used in the B-Model, we calculate all three indices and present the results in Table 1. Additionally, motivated from a recent work [asim_2025], we examine the ratios p/(2k1)p/(2k-1) and $p/g$ to understand how these derived quantities evolve with the $g$ and $k$, offering further insights into the internal consistency and comparative behavior of these inequality measures.
$R$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
1 0.983 0.957 0.973 1.011 0.973
3 0.971 0.926 0.956 1.016 0.954
6 0.951 0.886 0.933 1.023 0.931
11 0.919 0.827 0.902 1.028 0.900
16 0.886 0.775 0.876 1.029 0.874
21 0.852 0.726 0.853 1.030 0.852
30 0.796 0.653 0.818 1.027 0.822
35 0.762 0.617 0.801 1.025 0.809
40 0.731 0.583 0.785 1.022 0.798
46 0.695 0.546 0.768 1.019 0.786
51 0.665 0.517 0.754 1.016 0.778
$R$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
56 0.636 0.488 0.741 1.013 0.768
62 0.605 0.460 0.727 1.010 0.759
67 0.581 0.438 0.717 1.009 0.754
73 0.556 0.415 0.706 1.008 0.747
75 0.550 0.409 0.703 1.008 0.744
80 0.532 0.394 0.695 1.009 0.742
83 0.521 0.385 0.691 1.009 0.739
85 0.517 0.381 0.689 1.009 0.737
90 0.502 0.369 0.683 1.009 0.736
95 0.501 0.368 0.683 1.009 0.735
100 0.501 0.368 0.683 1.010 0.737
Table 1: B-Model with Exchange Range: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for B-Model with different range $R$ computed from the Eqn (5). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for B-model with different range ($R$).
Figure 2 provides a visual summary of how the inequality measures evolve as we vary the range parameter in the B-Model. In the upper-left panel, we can see the variation of $g$, $p$ and $k$ against $R$. In the upper-right panel, we illustrate the joint behaviors of the $g$-$k$ and $g$-$p$ pairs. As $R$ increases, $g$ and $k$ indices initially change in an almost perfectly linear manner with a slope $(3/8)$ as discussed in [bijin]. During this early phase, the values of $g$ and $k$ coincide at a point approximately $0.85$. Beyond this point, however, the relationship becomes distinctly nonlinear, indicating that the two measures respond differently to further increases in the range parameter. On the other hand as $R$ increases, $g$ and $p$ both increases together in similar manner and going to coincide at extreme inequality point $1$.
The lower-left panel of Figure 2 displays the variation of the ratio p/(2k1)p/(2k-1) with respect to $g$ and $k$. The pattern shows an initial increase of this ratio up to a certain threshold value; subsequently, as $g$ and $k$ continues to grow, the ratio begins to decline. Throughout the range, the ratio remains slightly above unity, reaching its maximum about a $3\%$ elevation near the point where $g$ and $k$ coincide (in this case, $g=k\sim 0.85$). This location has previously been identified as a precursor to criticality in several self-organized critical models [sand, manna, soumya]. The mild, non-monotonic rise toward this peak followed by a gradual decrease reflects a subtle shift in the balance between the $p$ and $k$ indices as the underlying wealth-exchange dynamics approach this near-critical region. A similar trend is observed in the lower-right panel, which shows how the ratio $p/g$ depends on $g$ and $k$ indices. In this case, the ratio increases sharply with increasing the indices $g$ and $k$, suggesting that $p$ grows at a faster rate relative to $g$.
Figure 2: B-Model with Exchange Range: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with range of exchanges ($R$) for B-Model (with total population, $N=100$ and ensemble averages over $10^{3}$ realizations). (a) Variation of $g$, $p$ and $k$ with range $R$; (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and the $g=k\approx 0.85$ point, also $g$ vs $p$-index curve; (c) Change of p/(2k1)p/(2k-1) with $g$ and $k$; (d) Change of $p/g$ with $g$ and $k$.
5.2 B-Model with Savings
In this case as well, we evaluate the same three inequality measures $g$, $p$ and $k$ to investigate how they behave and relate to one another when the model dynamics are controlled by the saving propensity $\lambda$ using Eqn. (3.1.2). Table 2 presents the computed values of $g$, $p$, and $k$ for a range of $\lambda$ in the B-Model using Eqns. (2), (3) and (4) respectively derived from Eqn. (1). Alongside these primary indicators, we also examine the derived ratios p/(2k1)p/(2k-1) and $p/g$, which provide additional insight into how the internal structure of inequality evolves as agents save a larger fraction of their wealth. By observing how these quantities vary with $g$ and $k$, we can better understand the sensitivity of the model’s inequality characteristics to changes in saving propensity.
$\lambda$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
0.001 0.983 0.957 0.974 1.010 0.974
0.01 0.983 0.957 0.973 1.011 0.973
0.03 0.983 0.955 0.972 1.011 0.972
0.05 0.982 0.951 0.970 1.012 0.970
0.07 0.982 0.950 0.970 1.012 0.968
0.09 0.981 0.949 0.969 1.012 0.968
0.10 0.981 0.949 0.969 1.011 0.967
0.20 0.979 0.943 0.966 1.013 0.964
0.30 0.975 0.936 0.961 1.014 0.959
0.40 0.972 0.928 0.957 1.015 0.954
0.50 0.968 0.918 0.952 1.016 0.948
0.60 0.962 0.905 0.945 1.018 0.941
0.70 0953 0.888 0.935 1.021 0.932
0.80 0.940 0.865 0.923 1.022 0.920
0.90 0.911 0.818 0.899 1.025 0.897
0.91 0.906 0.810 0.895 1.026 0.894
0.92 0.900 0.800 0.890 1.027 0.889
0.93 0.893 0.790 0.884 1.027 0.884
0.94 0.884 0.776 0.878 1.027 0.878
0.95 0.871 0.758 0.869 1.027 0.870
$\lambda$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
0.951 0.870 0.758 0.869 1.027 0.870
0.952 0.866 0.751 0.866 1.027 0.868
0.955 0.861 0.745 0.863 1.026 0.864
0.958 0.855 0.737 0.859 1.026 0.861
0.96 0.847 0.726 0.854 1.025 0.857
0.961 0.844 0.722 0.852 1.025 0.855
0.965 0.827 0.700 0.843 1.021 0.846
0.969 0.803 0.670 0.829 1.018 0.835
0.97 0.795 0.661 0.825 1.017 0.832
0.971 0.787 0.652 0.821 1.015 0.828
0.975 0.745 0.605 0.800 1.011 0.812
0.979 0.686 0.545 0.771 1.006 0.794
0.98 0.668 0.528 0.762 1.006 0.789
0.981 0.650 0.510 0.754 1.004 0.785
0.985 0.557 0.426 0.713 1.002 0.765
0.989 0.440 0.328 0.664 1.001 0.746
0.99 0.407 0.302 0.651 1.001 0.742
0.992 0.336 0.246 0.623 1.000 0.733
0.995 0.218 0.158 0.579 0.999 0.723
0.998 0.090 0.064 0.532 0.999 0.713
Table 2: B-Model with Savings: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for B-model with different saving propensities $\lambda$ computed from the Eqn. (3.1.2). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for B-model with different saving propensities $\lambda$.
In the upper-left panel of Figure 3, we present the relationship between $g$, $p$ and $k$ as the saving propensity $\lambda$ varies. The upper-right panel of the figure depicts the variation of $k$ and $p$ with respect to $g$. Similar to the earlier case with the range parameter, $g$ and $k$ indices exhibit an almost linear relationship in the initial regime with slope $(3/8)$. They coincide at approximately $g\approx k\approx 0.87$, after which the curves begin to diverge, indicating a departure from the initial linear dependence as savings become more dominant in the wealth-exchange dynamics. Further we can see that $p$ is also increasing with $g$ starting both from 0 and coincide at $1$.
The bottom-left panel of Figure 3 illustrates how the ratio p/(2k1)p/(2k-1) evolves with $g$ and $k$. Consistent with our previous observation, the ratio remains slightly above unity across all values of $\lambda$, reaching a maximum of about a $3\%$ elevation near the point where $g$ and $k$ coincide (in this case $g=k\approx 0.87$). For small values of $g$ and $k$, the ratio increases gradually up to this peak. However, as these indices approach their coincident value near $0.87$, the ratio drops sharply. This behavior underscores the sensitivity of the interplay between the $p$ and $k$ indices as the system approaches this near-critical regime. The bottom-right panel of the figure depicts the variation of the ratio $p/g$ as a function of $g$ and $k$. Here, a clear increasing trend emerges: as $g$ and $k$ increases, the ratio $p/g$ consistently increases, following the pattern illustrated in the plot. This monotonic increase indicates that $p$ grows more rapidly relative to $g$.
Figure 3: B-Model with Savings: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with saving propensity ($\lambda$) for B-Model (with total population, $N=100$ and ensemble averages over $10^{3}$ realizations). (a) Variation of $g$, $p$ and $k$ with savings $\lambda$; (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and the $g=k\approx 0.87$ point, also $g$ vs $p$-index curve; (c) Change of p/(2k1)p/(2k-1) with $g$ and $k$; (d) Change of $p/g$ with $g$ and $k$.
5.3 C-Model with Savings
In this scenario as well, we evaluate the same set of three inequality indicators to understand how inequality evolves within the C-Model in presence of saving propensity, $\lambda$ following Eqn. (7). Table 3 reports the numerical values of these indices for various choices of $\lambda$ using Eqns. (1)–(4). Along with the primary measures, we also examine the derived ratios p/(2k1)p/(2k-1) and $p/g$, which provide additional insight into the structural changes in inequality as the saving propensity is varied. By tracking how these ratios respond to changes in $g$ and $k$, we gain a clearer picture of the internal consistency and comparative behavior of the inequality measures in the C-Model.
$\lambda$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
0.01 0.990 0.983 0.988 1.007 0.993
0.10 0.989 0.979 0.986 1.008 0.990
0.14 0.989 0.977 0.984 1.008 0.988
0.18 0.989 0.975 0.984 1.008 0.986
0.20 0.989 0.974 0.983 1.007 0.985
0.23 0.988 0.973 0.983 1.008 0.985
0.27 0.988 0.971 0.982 1.007 0.983
0.30 0.987 0.968 0.980 1.009 0.981
0.31 0.987 0.967 0.979 1.010 0.980
0.35 0.986 0.962 0.976 1.010 0.976
0.39 0.985 0.959 0.974 1.011 0.974
0.40 0.984 0.959 0.975 1.009 0.974
0.44 0.983 0.954 0.971 1.011 0.970
0.48 0.981 0.947 0.967 1.012 0.966
0.50 0.980 0.945 0.966 1.013 0.965
0.52 0.978 0.940 0.964 1.013 0.961
0.56 0.975 0.931 0.960 1.013 0.955
0.60 0.971 0.923 0.954 1.016 0.950
0.61 0.971 0.922 0.955 1.013 0.949
0.65 0.965 0.907 0.947 1.015 0.939
0.69 0.956 0.889 0.937 1.018 0.929
0.70 0.953 0.882 0.933 1.018 0.925
0.73 0.944 0.864 0.925 1.018 0.916
$\lambda$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
0.77 0.925 0.829 0.906 1.020 0.897
0.80 0.908 0.801 0.892 1.020 0.882
0.81 0.899 0.786 0.885 1.021 0.875
0.815 0.892 0.778 0.881 1.022 0.872
0.82 0.890 0.774 0.879 1.021 0.870
0.83 0.880 0.758 0.872 1.019 0.862
0.84 0.870 0.745 0.864 1.022 0.857
0.85 0.857 0.729 0.857 1.020 0.850
0.857 0.843 0.709 0.848 1.019 0.841
0.86 0.837 0.704 0.844 1.023 0.840
0.87 0.821 0.681 0.834 1.021 0.830
0.88 0.798 0.655 0.821 1.020 0.820
0.89 0.773 0.627 0.808 1.019 0.812
0.90 0.744 0.595 0.792 1.018 0.801
0.911 0.708 0.561 0.775 1.020 0.792
0.922 0.660 0.513 0.752 1.017 0.778
0.933 0.608 0.465 0.729 1.016 0.764
0.944 0.544 0.409 0.702 1.013 0.753
0.955 0.471 0.349 0.673 1.010 0.741
0.966 0.383 0.281 0.639 1.009 0.735
0.977 0.280 0.203 0.601 1.005 0.726
0.988 0.157 0.112 0.556 1.002 0.713
0.999 0.015 0.010 0.505 1 0.710
Table 3: C-Model with Savings: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for C-model with different saving propensities $\lambda$ computed from the Eqn. (7). Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for C-model with different saving propensities $\lambda$.
The upper-left panel of Figure 4 illustrates how $g$, $p$ and $k$ co-varies with the saving propensity $\lambda$ in the C-Model or Yard-Sale Model framework. Further in the upper-right panel of this figure, we observe that at low and intermediate values of $g$, $k$ evolve in near synchrony, producing an almost linear trend. They intersect at a point close to $g=k\approx 0.86$. Beyond this intersection, however, the relation begins to bend, signaling a clear departure from linearity as higher saving propensities increasingly influence the wealth-exchange dynamics. Here we can also see the variation of $p$ with $g$, these two measures coincide only at 0 and $1$.
The lower-left panel shows the behavior of the ratio p/(2k1)p/(2k-1) as a function of $g$ as well as $k$. The ratio remains slightly above unity, reaching a modest peak—about a $3\%$ increase near the point where $g$ and $k$ coincide. For small values of $g$ and $k$, the ratio increases gradually before attaining this peak. As the system approaches the coincident value (here $g=k\approx 0.86$), however, the ratio undergoes a pronounced decline, indicating a sharp reduction in the disparity between the $p$ and $k$ indices. The lower-right panel displays the ratio $p/g$ across $g$ and $k$. Here, a clear monotonic increase is observed: as $g$ and $k$ increases, the value of $p/g$ steadily increases. This pattern indicates that $p$ becomes progressively larger relative to $g$ consistent with the trend captured in the plot.
Figure 4: C-Model with Savings: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with saving propensity ($\lambda$) for C-Model (with total population, $N=100$ and ensemble averages over $10^{3}$ realizations). (a) Variation of $g$, $p$ and $k$ with saving propensity $\lambda$; (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and the $g=k\approx 0.86$ point, also $g$ vs $p$-index curve; (c) Change of p/(2k1)p/(2k-1) with $g$ and $k$; (d) Change of $p/g$ with $g$ and $k$.
5.4 P-Model
We simulated income distributions governed by the Pareto law following Eqns. (8) and (9) and evaluated the associated inequality measures- namely, the Gini coefficient $g$, the Pietra index $p$, and the Kolkata index $k$ using Eqns. (1)–(4). The minimum income is fixed at unity, and all three indices exhibit a systematic and monotonic dependence on the Pareto shape parameter $\alpha$. As $\alpha$ approaches its lower bound of $1^{+}$, the distribution enters the regime of maximal permissible inequality while still maintaining a finite mean. Correspondingly, the indices converge toward their characteristic limiting values of approximately $0.85$.
A decrease in $\alpha$ enhances the heaviness of the power-law tail, implying an increasingly concentrated allocation of income among a diminishing fraction of individuals. This redistribution toward the upper tail is quantitatively captured by the pronounced rise in both $g$ and $k$ indices, each of which serves as a robust scalar measure of inequality. Their simultaneous escalation with decreasing $\alpha$ demonstrates that the broadening of the tail is directly reflected in the macroscopic inequality descriptors of the system.
$\alpha$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
1.001 0.875 0.765 0.872 1.028 0.875
1.243 0.655 0.524 0.750 1.047 0.799
1.545 0.476 0.365 0.673 1.055 0.766
1.847 0.371 0.280 0.632 1.059 0.754
2.149 0.304 0.227 0.607 1.059 0.747
2.209 0.293 0.219 0.603 1.060 0.747
2.451 0.256 0.590 0.190 1.060 0.743
2.752 0.222 0.165 0.578 1.061 0.742
3.054 0.196 0.145 0.568 1.061 0.740
3.356 0.175 0.129 0.561 1.060 0.740
3.658 0.158 0.117 0.555 1.061 0.739
3.960 0.144 0.107 0.550 1.061 0.738
4.262 0.133 0.098 0.546 1.062 0.738
4.564 0.123 0.091 0.543 1.062 0.738
4.866 0.115 0.084 0.540 1.062 0.737
5.168 0.107 0.079 0.537 1.062 0.738
$\alpha$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
5.470 0.101 0.074 0.535 1.063 0.738
5.772 0.095 0.070 0.533 1.061 0.738
6.074 0.090 0.066 0.531 1.063 0.737
6.376 0.085 0.063 0.529 1.063 0.737
6.678 0.081 0.060 0.528 1.062 0.737
6.980 0.077 0.057 0.527 1.062 0.737
7.282 0.074 0.054 0.526 1.060 0.736
7.584 0.071 0.052 0.524 1.061 0.736
7.886 0.068 0.045 0.523 1.060 0.736
8.188 0.065 0.048 0.523 1.060 0.736
8.490 0.063 0.046 0.522 1.062 0.737
8.792 0.060 0.044 0.521 1.062 0.736
9.094 0.058 0.043 0.520 1.062 0.737
9.396 0.056 0.041 0.519 1.067 0.737
9.698 0.054 0.040 0.519 1.064 0.737
10.000 0.053 0.039 0.518 1.063 0.736
Table 4: P-Model: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for P-model with different exponent $\alpha$, calculated using Eqn. (9), along with the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for P-model with different exponent $\alpha$.
The upper-left panel of Figure 5 shows how the inequality measures $g$, $p$, and $k$ evolve with the exponent $\alpha$. In the upper right panel we can see that $k$ and $g$ mostly follow a linear relation, till they both become equal at $\approx 0.88$, at $\alpha=1.001$. We also see the $p$ is always lesser than $g$ for the values of $\alpha$ considered.
The lower-left panel shows the behavior of the ratio p/(2k1)p/(2k-1) as a function of $g$ as well as $k$. The ratio is always greater than 1, with its value decreasing from $\approx 1.07$ to $\approx 1.03$. The lower-right panel displays the ratio $p/g$ across $g$ and $k$. A clear monotonic increase is observed.
Refer to captionRefer to caption
Figure 5: P-Model: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with ($\alpha$) for P-Model for minimum income=11. (a) Variation of $g$, $p$ and $k$ with exponent $\alpha$; (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and the $g=k\approx 0.88$ point, also $g$ vs $p$-index curve; (c) Change of p/(2k1)p/(2k-1) with $g$ and $k$; (d) Change of $p/g$ with $g$ and $k$.
5.5 CS-Model
The Kolkata–Gini relation obtained from the sequence of Cantor generations, initially exhibits small but systematic oscillations around an otherwise smooth and approximately monotonic curve. Here too, we calculate $g$, $k$ and $p$, along with the ratios p/(2k1)p/(2k-1) and $p/g$, to understand how inequality evolves with each fractal generation. The table 4 shows values at 5-generation intervals from generation 5 onward, with generation 4 included separately. .
$n$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
4 0.359 0.634 0.280 1.041 0.781
5 0.400 0.290 0.641 1.026 0.725
10 0.548 0.409 0.700 1.021 0.748
15 0.645 0.493 0.743 1.016 0.765
20 0.715 0.558 0.776 1.011 0.779
25 0.769 0.610 0.803 1.007 0.793
30 0.811 0.653 0.825 1.004 0.805
35 0.844 0.690 0.845 1.000 0.817
40 0.871 0.721 0.860 1.002 0.828
45 0.893 0.749 0.873 1.003 0.839
50 0.911 0.773 0.885 1.003 0.849
55 0.925 0.794 0.896 1.002 0.858
60 0.937 0.812 0.906 1.001 0.867
65 0.947 0.829 0.914 1.001 0.875
70 0.956 0.845 0.922 1.001 0.884
75 0.963 0.860 0.929 1.001 0.893
80 0.968 0.873 0.936 1.001 0.901
85 0.973 0.884 0.942 1.001 0.908
90 0.977 0.895 0.947 1.001 0.915
95 0.981 0.904 0.952 1.001 0.921
100 0.984 0.912 0.956 1.000 0.927
$n$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
105 0.986 0.920 0.960 1.000 0.933
110 0.988 0.926 0.963 1.000 0.938
115 0.990 0.933 0.966 1.001 0.942
120 0.992 0.938 0.969 1.000 0.946
125 0.993 0.943 0.971 1.000 0.950
130 0.994 0.948 0.974 1.000 0.953
135 0.995 0.952 0.976 1.000 0.957
140 0.996 0.956 0.978 1.000 0.960
145 0.996 0.960 0.980 1.000 0.963
150 0.997 0.963 0.981 1.000 0.966
155 0.997 0.966 0.983 1.000 0.969
160 0.998 0.969 0.984 1.000 0.971
165 0.998 0.972 0.9860 1.000 0.973
170 0.998 0.974 0.987 1.000 0.976
175 0.998 0.976 0.988 1.000 0.977
180 0.999 0.978 0.989 1.000 0.979
185 0.999 0.980 0.990 1.000 0.981
190 0.999 0.981 0.991 1.000 0.982
195 0.999 0.982 0.991 1.000 0.983
199 0.999 0.984 0.992 1.000 0.984
200 0.999 0.984 0.992 1.000 0.985
Table 5: CS-Model: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for CS-model with different generations ($n$), where the overlap magnitude $Y_{n}(t)$ is calculated using Eqn. (10) along with the computed ratios p/(2k1)p/(2k-1) and $p/g$ (motivated by [asim_2025]).
The upper-left panel of Figure 6 shows how the inequality measures $g$, $p$, and $k$ evolve with the generation number in the CS model. The upper-right panel further reveals that, for low to intermediate values of $g$, the Kolkata index $k$ tracks $g$ closely, resulting in an almost linear movement. The two curves intersect near $g=k\approx 0.84$. Beyond this point, the relationship becomes non linear. The panel also displays how the Pietra index $p$ varies with $g$; these two measures coincide only at the maximal inequality value $p=g=1$.
The lower-left panel shows the behavior of the ratio p/(2k1)p/(2k-1) as a function of $g$ as well as $k$. The ratio is always greater than 1, with its value decreasing from $\approx 10\%$ until it reaches 1 for the first time when $g=k\approx 0.84$. The lower-right panel displays the ratio $p/g$ across $g$ and $k$. Here too, a clear monotonic increase is observed.
Refer to captionRefer to caption
Figure 6: CS-Model: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with $n\in[4,200]$ for CS-Model. (a) Time ($t$) series of overlap magnitude $Y_{4}(t)$ of the 4th generation of the cantor set; (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and the $g=k\approx 0.84$ point, also $g$ vs $p$-index curve; (c) Change of p/(2k1)p/(2k-1) with $g$ and $k$; (d) Change of $p/g$ with $g$ and $k$.
5.6 BK-Model
We analyze the temporal evolution of the elastic energy released and the corresponding kkgg and ppgg relations. For the BK-Model, the $k$ versus $g$ plot exhibits an approximately linear trend with slope $\simeq 0.375$, suggesting a nontrivial correspondence with the Lorenz-curve geometry observed in the P-model, despite the fundamentally different underlying dynamics. In the BK-Model, the inequality measures quantify the instantaneous distribution of spring forces across the block ensemble, providing a time-resolved characterization of stress heterogeneity during loading and slip events.
Refer to captionRefer to caption
Figure 7: BK-model: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with number of blocks (NN) = $100$, $\alpha=1000$, $k_{c}/k_{p}=1$ and $\sigma=0.01$ for BK-Model. (a) Elastic energy released with respect to time ($t$); (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and also $g$ vs $p$-index curve; (c) pp/(2k1)(2k-1) vs. $g$ and $k$ and; (d) pp/$g$ vs. $g$ and $k$ plotted beyond the transient phase. Here the equation of motion of the blocks and the frictional force of them are given by Eqns. (11) and (12) respectively.
In the top right panel, we plot $k$ and $p$ with respect to $g$. Surprisingly, in this regime of the model, $k$ and $g$ closely follow a linear relationship with slope $3/8$, similar to the previous examples.
In the bottom left panel we show p/(2k1)p/(2k-1), where the values drop below 1, ranging from $\approx 0.98$ to $\approx 1.08$İn the bottom right panel, we plot $p/g$ after the initial transient period (first 100 steps) has passed and a steady state is reached, since the system starts. In the Burridge-Knopoff model, the initial evolution is dominated by stress-loading transients arising from the imposed initial conditions. During this phase, event statistics and associated inequality measures such as the Gini and Kolkata indices exhibit strong time dependence and are not representative of the intrinsic dynamics. In the bottom right panel, the ratio $p/g$ ranges from $\approx 0.69$ and to $\approx 0.82$
6 Summery and Concluding remarks
In this study, we have performed a comparative numerical investigation of inequality measures (namely of Gini [gini], Pietra [pietra] and Kolkata [k-index1, k-index3] indices) across a broad spectrum of model systems: they range from two kinetic wealth-exchange models (namely the Banerjee Model [B_Model, B_Model1] and Chakraborti model [C_Model], more popularly known as Yard-Sale model [YS, YS1]), to synthetic Pareto model [Pareto, Pareto1] and two earthquake-generating models (namely the Chakrabatrti-Stinchcombe model [CS_Model, CS1, pbhattacharya] and the Burridge-Knopoff model [langer_1994, BK_Model]). Using the Lorenz curves (see Fig. 1), generated numerically for each of these model systems for different parameter values (using Eqns. (5),(3.1.2) for B-Model, Eqn. (7) for C-Model, Eqn. (9) for P-Model, Eqn. (10) for CS-Model and Eqns. (11),(12) for the BK-Model), we estimate the different inequality indices, namely the Gini index ($g$), Pietra index ($p$), and Kolkata index ($k$). The results of our numerical analysis for the above mentioned inequality indices are given and shown respectively in Tables 1-5 and Figs. 2-7. It may be noted that while switching from the econophysical wealth exchange models (B and C models) which have got dominantly exponential (to Gamma-like) distributions and the two earthquake simulating models (CS and BK models) which have got (Guttenberg-Richter like) dominantly power law distributions, we considered and compared also the results for the perfect power-law Pareto distribution as well.
We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models, the two-fractal model and the Pareto distributions) remain a little above unity, which was predicted to be exactly equal to unity [asim_2025]. The value of this quantity for the two socio-economic (B and C models) and the earthquake simulating CS model is seen to deviate a little above unity, bounded by a maximum of $4\%$ near $g=k=0.86\pm 0.02$ which was identified earlier [manna, soumya] to be the precursor point of criticality in several self-organized critical models. For the Pareto distribution, the ratio p/(2k1)p/(2k-1) deviates slightly above unity, reaching a maximum of about $7\%$ (see Fig. 5(c)). In the BK-Model, however, the value of p/(2k1)p/(2k-1) even drops below unity in certain regimes (see Fig. 7(c)). The predicted values (assuming a simple polynomial form for the Lorenz function [asim_2025]) of $p/g=3/4$ seems to deviate quite a bit (in the range $0.69$ to about $1.0$; see Tables 1-5), while the existence of the relation $k=1/2+(3/8)g$ [asim_2025] seems to be valid for all the models considered here in the region of lower $g$ values (see Figs. 2(d) - 7(d)).
These observations of quantitatively similar behaviors for the inequality indices, studied here for different socio-economic and geophysical models, may provide some useful and coherent comparative framework for studying the universal features of the statistics emerging from across disparate dynamical systems.

### v2

@@
 Abstract
-We present a numerical study of several inequality measures across two kinetic wealth-exchange models with extreme inequality features (namely the Banerjee model, and the Chakraborti or Yard-Sale model) and two earthquake simulating models (namely the Chakrabarti-Stinchcombe two-fractal overlap model and the nonlinear dynamical Burridge-Knopoff model), and a synthetic Pareto distribution.
+We present a numerical study of several inequality measures across two kinetic wealth-exchange models with extreme inequality features (namely the Banerjee model, and the Chakraborti or Yard-Sale model) and two earthquake simulating models (namely the Chakrabarti-Stinchcombe two-fractal overlap model and the nonlinear dynamical Burridge-Knopoff model).
 For each model we compute numerically the Lorenz function for the respective models’ wealth, overlap magnitude or avalanche distributions.
@@
 We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models and the two-fractal overlap model) remain a little above unity (theoretically predicted value) and deviating a little higher by a maximum of $4\%$ near $g=k\simeq 0.86$, which was identified earlier to be the precursor point of criticality in several self-organized critical models ($k=0.80$ corresponds to Pareto’s 80-20 law).
-In the Burridge-Knopoff model for some instances of time, the value of p/(2k1)p/(2k-1) drops a little below unity.
+The Burridge–Knopoff model shows larger deviations from the $k$-$g$ relation, but does not access the high-inequality regime where an intersection occurs.
 This and some other quantitatively similar behaviors of the inequality indices across socio-economic and geophysical models may provide a coherent and comparative framework for identifying the subtle features in the statistics of such disparate dynamical systems.
@@
 […]
+Inequality-how unevenly outcomes are distributed, shows up not only in economics, but also in many driven physical systems where typically fewer larger events (avalanches) tend to dominate the major scaling behavior.
+In this work, we use a common quantitative language to compare ‘extreme inequality’ across two kinetic wealth-exchange models and two earthquake-inspired models.
+From the simulated distributions (wealth, fractal-overlap magnitudes, and slip avalanches), we construct the Lorenz curve and extract three standard inequality measures: the Gini index ($g$), the Pietra index ($p$), and the Kolkata index ($k$).
+By systematically varying key control parameters in each model (saving propensity in the economic models; system size or generation/block number in the geophysical models), we show that a specific composite ratio, p/(2k1)p/(2k-1), stays close to its theoretical benchmark—typically slightly above unity and deviating by at most $\sim 8\%$, generally near a previously identified precursor or threshold regime around $g=k\approx 0.86$.
+Overall, these cross-model regularities suggest that inequality indices can serve as sensitive, comparable diagnostics for recognizing subtle changes in the statistics of disparate complex systems, from markets to fault dynamics.
 1 Introduction
@@
 To compare these behaviors within a unified framework, we numerically quantify inequality in four classes of models: two kinetic wealth-exchange models, a synthetic Pareto distribution, and two earthquake-simulating models.
-For the socio-economic domain, early models introduced by Dragulescu and Yakovenko [DY_2000] and Chakraborti and Chakrabarti [CC_2000] demonstrated that simple stochastic exchange rules can generate realistic wealth distributions.
-Extensions involving fixed or distributed saving propensity [CCM_2004], interaction heterogeneity, and asymmetric exchange have yielded richer phenomenology.
-The Chakraborti Model [C_Model] or also known as Yard-Sale model [YS], later extended by Boghosian and collaborators [YS1], captures multiplicative dynamics and extreme wealth condensation.
-The Banerjee Model [B_Model, B_Model1] represents another natural kinetic exchange framework, whose intrinsic trading dynamics lead to steady-state configurations characterized by extreme concentration of wealth among a small subset of agents.
+For the socio-economic domain, early models introduced by Dragulescu and Yakovenko [16] and Chakraborti and Chakrabarti [11] demonstrated that simple stochastic exchange rules can generate realistic wealth distributions.
+Extensions involving fixed or distributed saving propensity [13], interaction heterogeneity, and asymmetric exchange have yielded richer phenomenology.
+The Chakraborti Model [12] or also known as Yard-Sale model [8], later extended by Boghosian and collaborators [7], captures multiplicative dynamics and extreme wealth condensation.
+The Banerjee Model [5, 18] represents another natural kinetic exchange framework, whose intrinsic trading dynamics lead to steady-state configurations characterized by extreme concentration of wealth among a small subset of agents.
 In this study, we focus on the Banerjee Model, where inequality is tuned by varying the range of exchange and the saving propensity, and the Chakraborti Model (or Yard-Sale model), where saving propensity acts as the central control parameter.
-To broaden the scope beyond socio-economic systems, we also examine inequality associated with Pareto-type distributions [Pareto, Pareto1] by tuning their shape parameter $\alpha$, given the ubiquity of power laws in systems exhibiting self-organized criticality.
-Since such distributions are central to geophysical phenomena, we analyze inequality patterns in two different earthquake models, the Chakrabarti–Stinchcombe two-fractal overlap model [CS_Model, CS1], where event sizes arise from Cantor-set overlaps, and the nonlinear Burridge–Knopoff spring-block model [langer_1994, BK_Model], where block–spring dynamics generate heterogeneous slip events.
-Quantifying inequality across these systems , we therefore investigate three Lorenz curve [lor, lor1] based measures: the Gini index (gg[gini], captures the average deviation from perfect equality; The Pietra (or Hoover) index (pp[pietra, pietra1], measures the maximum imbalance in cumulative wealth and the more recently introduced Kolkata index (kk[k-index, k-index1, k-index2, k-index3], identifies the wealth share of the richest fraction of the population and has found applications in diverse socio-economic contexts.
-For each system, we systematically vary the relevant control parameters such as the exchange range ($R$) and saving propensity ($\lambda$) in the Banerjee Model, saving propensity ($\lambda$) in the Chakraborti Model, the exponent ($\alpha$) in the synthetic Pareto distribution, the Cantor generation index in Chakrabarti–Stinchcombe two-fractal overlap model, and the time evolution of the Burridge–Knopoff model and compute the corresponding inequality indices.
-In a recent work [asim_2025] authors studied the ratios p/(2k1)p/(2k-1) and $p/g$ in some financial systems.
+To broaden the scope beyond socio-economic systems, we also examine inequality associated with Pareto-type distributions [28, 26] by tuning their shape parameter $\alpha$, given the ubiquity of power laws in systems exhibiting self-organized criticality.
+Since such distributions are central to geophysical phenomena, we analyze inequality patterns in two different earthquake models, the Chakrabarti Stinchcombe two-fractal overlap model [10, 23], where event sizes arise from Cantor-set overlaps, and the nonlinear Burridge Knopoff spring-block model [9, 27], where block–spring dynamics generate heterogeneous slip events.
+Quantifying inequality across these systems , we therefore investigate three Lorenz curve [24, 1] based measures: the Gini index (gg[21], captures the average deviation from perfect equality; The Pietra (or Hoover) index (pp[29, 17], measures the maximum imbalance in cumulative wealth and the more recently introduced Kolkata index (kk[20, 14, 4, 3], identifies the wealth share of the richest fraction of the population and has found applications in diverse socio-economic contexts.
+For each system, we systematically vary the relevant control parameters such as the exchange range ($R$) and saving propensity ($\lambda$) in the Banerjee Model, saving propensity ($\lambda$) in the Chakraborti Model, the exponent ($\alpha$) in the synthetic Pareto distribution, the Cantor generation index in Chakrabarti Stinchcombe two-fractal overlap model, and the time evolution of the Burridge Knopoff model and compute the corresponding inequality indices.
+In a recent work [19] authors studied the ratios p/(2k1)p/(2k-1) and $p/g$ in some financial systems.
 We further examine the ratios p/(2k1)p/(2k-1), $p/g$ for these kinetic exchange models, the two earthquake simulating models, and the Pareto distribution to investigate the variation of these and other relationships among the inequality measures $g$, $p$ and $k$.
@@
 Quantifying inequality in any distribution whether describing wealth, event sizes, or released energy (hereafter called wealth) requires statistical measures that capture heterogeneity in a systematic manner.
-In this study, we numerically examine three statistical measures based on the Lorenz function [lor, lor1], namely the Gini index [gini], the Pietra (or Hoover) index [pietra, pietra1], and the Kolkata index [k-index, k-index2].
+In this study, we numerically examine three statistical measures based on the Lorenz function [24, 1], namely the Gini index [21], the Pietra (or Hoover) index [29, 17], and the Kolkata index [20, 4].
 The Lorenz function (or Lorenz curve) $L(x)$ is a foundational tool for visualizing and quantifying inequality.
@@
 $L(x)=\frac{\displaystyle\int_{0}^{x}y\,P(y)\,dy}{\displaystyle\int_{0}^{\infty}y\,P(y)\,dy},\qquad 0\leq x\leq 1.$ (1)
-Here, $x$ denotes the bottom $x$ fraction of the population, and $L(x)$ gives the corresponding fraction of the wealth they hold.
+Here, $x$ denotes the bottom $x$ fraction of the population, and $L(x)$ gives the corresponding cumulative fraction of the wealth they hold.
 For a perfectly equal distribution, $L(x)=x$, producing a 45-degree line known as the perfect equality line (see in Fig.
@@
 The $k$-index varies from $1/2$ (perfect equality) to $1$ (perfect inequality).
-In many empirical systems and theoretical models, $k$ displays near-universal scaling behaviour in relation to $g$ or $p$, which motivates our numerically comparative study across two kinetic wealth exchange models and two earthquake simulating models [sand, manna, soumya, asim_2025].
+In many empirical systems and theoretical models, $k$ displays near-universal scaling behaviour in relation to $g$ or $p$, which motivates our numerically comparative study across two kinetic wealth exchange models and two earthquake simulating models [2, 25, 15, 19].
 3 Models
@@
 3.1 Kinetic Wealth-Exchange Models
-Kinetic wealth exchange models given by Dragulescu and Yakovenko [DY_2000] describes wealth evolution through pairwise stochastic exchanges among agents and Chakraborti and Chakrabarti through saving propensities [CC_2000].
+Kinetic wealth exchange models given by Dragulescu and Yakovenko [16] describes wealth evolution through pairwise stochastic exchanges among agents and Chakraborti and Chakrabarti through saving propensities [11].
 Below we describes our considered two kinetic wealth exchange models.
 3.1.1 B-Model with Exchange Range
-The Banerjee model [B_Model, B_Model1] (or B-Model in short hereafter) introduces a constrained exchange rule in which each agent interacts only within a selected exchange range $R$ in the wealth-ordered list.
+The Banerjee model [5, 18] (or B-Model in short hereafter) introduces a constrained exchange rule in which each agent interacts only within a selected exchange range $R$ in the wealth-ordered list.
 At each time step $t$:
@@
 where $\epsilon$ is a random number drawn uniformly from $(0,1)$.
-This model naturally generates strong wealth condensation for small $R$, producing highly unequal steady states—though not to the extreme limit observed in the C-Model (discussed later), where all wealth collapses into a single agent.
+This model naturally generates strong wealth condensation for small $R$, producing highly unequal steady states, though not to the extreme limit observed in the C-Model (discussed later), where all wealth collapses into a single agent.
 For each choice of $R$, we compute the steady-state Lorenz geometry and associated inequality indices after ensemble averaging.
@@
 3.1.3 C-Model with Savings
-The Chakraborti Model [C_Model] (or C-Model in short hereafter) also widely known as Yard-Sale model [YS] is based on multiplicative dynamics in which the amount at stake depends on the wealth of the poorer agent.
+The Chakraborti Model [12] (or C-Model in short hereafter) also widely known as Yard-Sale model [8] is based on multiplicative dynamics in which the amount at stake depends on the wealth of the poorer agent.
 With saving propensity $\lambda\in[0,1)$, the effective tradable wealth becomes $(1-\lambda)\min(w_{i},w_{j})$.
@@
 3.2 P-Model
-We investigate inequality in synthetic Pareto-type distribution [Pareto, Pareto1] (or P-Model in short hereafter) characterized by the probability density function,
+We investigate inequality in synthetic Pareto-type distribution [28, 26] (or P-Model in short hereafter) characterized by the probability density function,
 $p(x)=\frac{\alpha x_{m}^{\alpha}}{x^{\alpha+1}},\qquad x\geq x_{m},$ (8)
@@
 $L(p)=1-(1-p)^{\frac{\alpha-1}{\alpha}},$ (9)
-where $p$ denotes the cumulative fraction of the population, and $L(p)$ is the corresponding fraction of total wealth held by that population.
+where $p$ denotes the fraction of the population, and $L(p)$ is the corresponding cumulative fraction of total wealth.
 For each $\alpha$ we generate large synthetic datasets, compute the Lorenz curve, and evaluate the associated inequality indices.
@@
 We now move to two classes of geophysical models that simulate earthquake like event size statistics.
-The Chakrabarti-Stinchcombe fractal overlap construction generates multifractal, heavy-tailed spatial measures that produce inequality through time variation of the overlap magnitude, whereas the Burridge–Knopoff model gives rise to inequality in terms of spatially varying spring force.
+The Chakrabarti Stinchcombe fractal overlap construction generates multifractal, heavy-tailed spatial measures that produce inequality through time variation of the overlap magnitude, whereas the Burridge Knopoff model gives rise to inequality in terms of spatially varying spring force.
 Both models probe different but complementary aspects of earthquake dynamics, namely fault geometry and stress distribution during an earthquake.
-The Chakrabarti–Stinchcombe model exhibits a power-law scaling in the cumulative distribution of overlap magnitudes in the asymptotic limit of large generation number and large overlaps, arising purely from the construction of the underlying fractlas [pbhattacharya].
-In contrast, the Burridge-Knopoff model displays power-law event-size statistics in specific dynamical regimes, where nonlinear friction and collective interactions drive the system toward scale-free behavior.
-Thus, although both models can produce similar phenomenological power-law statistics, the underlying mechanisms are fundamentally different– geometric hierarchy in the former and dynamical evolution in the latter.
+The Chakrabarti Stinchcombe model exhibits a power-law scaling in the cumulative distribution of overlap magnitudes in the asymptotic limit of large generation number and large overlaps, arising purely from the construction of the underlying fractals [6].
+In contrast, the Burridge Knopoff model displays power-law event-size statistics in specific dynamical regimes, where nonlinear friction and collective interactions drive the system toward scale-free behavior.
+Thus, although both models can produce similar phenomenological power-law statistics, the underlying mechanisms are fundamentally different, geometric hierarchy in the former and dynamical evolution in the latter.
 3.3.1 CS-Model
-The Chakrabarti Stinchcombe model [CS_Model, CS1, pbhattacharya] (CS-Model in short hereafter) based on the overlap of fractal fault surfaces (modeled using Cantor sets), focuses on fault geometry and the distribution of overlaps that correspond to stored elastic energy.
+The Chakrabarti Stinchcombe model [10, 23, 6] (CS-Model in short hereafter) based on the overlap of fractal fault surfaces (modeled using Cantor sets), focuses on fault geometry and the distribution of overlaps that correspond to stored elastic energy.
 It is therefore a model of geometric heterogeneity: it explains why slip events follow broad distributions (e.g., power laws) purely from the self-similar structure of the fault interface suitable for Lorenz-curve–based inequality analysis.
-Earthquake magnitudes empirically follow the Gutenberg–Richter law, $\log_{10}N_{r}(M>M_{0})=a-bM$, with $N_{r}(M>M_{0})$ denoting the number of events above magnitude $M_{0}$, and $b\simeq 1$.
+Earthquake magnitudes empirically follow the Gutenberg–Richter law,
+$\log_{10}N_{r}(M>M_{0})=a-bM,$
+with $N_{r}(M>M_{0})$ denoting the number of events above magnitude $M_{0}$, and $b\simeq 1$.
 The model captures this scale-free behavior by treating a fault as two identical self-affine fractals, typically $n^{th}$-generation Cantor sets with $2^{n}$ segments of length $3^{-n}$, sliding over each other.
@@
 $N_{r}\!\left(2^{\,n-k}\right)=2^{k}\binom{n}{k}$ (10)
-which normalize to the binomial probabilities $\Pr(2^{\,n-k})=\binom{n}{k}\left(\tfrac{1}{3}\right)^{n-k}\left(\tfrac{2}{3}\right)^{k}$; the most probable overlap scales as $\hat{Y}_{n}\sim(2^{n})^{1/3}$.
+which normalize to the binomial probabilities
+$\Pr(2^{\,n-k})=\binom{n}{k}\left(\tfrac{1}{3}\right)^{n-k}\left(\tfrac{2}{3}\right)^{k};$
+the most probable overlap scales as $\hat{Y}_{n}\sim(2^{n})^{1/3}$.
 CS-Model captures the static, structural statistics of fault roughness providing insight into how geometric irregularity shape the statistics of earthquakes.
 3.3.2 BK-Model
-In contrast, the Burridge–Knopoff Model [langer_1994, BK_Model] (BK-Model in short hereafter) is a dynamical and mechanical model of earthquake generation, where blocks connected by springs and subject to friction evolve over time.
+In contrast, the Burridge Knopoff Model [9, 27] (BK-Model in short hereafter) is a dynamical and mechanical model of earthquake generation, where blocks connected by springs and subject to friction evolve over time.
 It is used to study stick–slip dynamics and stress redistribution.
@@
 where $F_{s}$ is the static friction force $\mu$ is related to the friction weakening rate, and $\sigma$ is related to the friction drop parameter, when it changes from static to kinetic.
-Using the stick–slip friction law—with parameters $\alpha$ (friction weakening rate) and $\sigma$ (friction drop), the model can reproduce Gutenberg–Richter–like magnitude statistics, exhibit chaotic dynamics despite being deterministic, and transition between periodic, chaotic, and scale-free regimes depending on parameters.
+Using the stick–slip friction law—with parameters $\delta$ (friction weakening rate) and $\sigma$ (friction drop), the model can reproduce Gutenberg–Richter–like magnitude statistics, exhibit chaotic dynamics despite being deterministic, and transition between periodic, chaotic, and scale-free regimes depending on parameters.
 Although more sophisticated rate-and-state (R&S) friction laws exist, we use the stick–slip rule which is computationally simpler.
@@
 (10), with $n$ ranging from 4 to 200.
-The inequality measures $g$, $k$, and $p$ are computed from the corresponding Lorenz curves, where the overlap magnitude is plotted along the abscissa and the cumulative frequency along the ordinate.
+The inequality measures $g$, $k$, and $p$ are computed from the corresponding Lorenz curves, where the cumulative normalized overlap magnitude is plotted along the ordinate against the cumulative fraction of events along the abscissa.
 These measures provide a compact quantitative characterization of the degree of heterogeneity in the spatial overlap and enable us to track the evolution of inequality as the fractal resolution is progressively increased.
-For the BK model, we perform numerical simulations to investigate the emergence of collective slip dynamics for a given set of physical parameters.
+For the BK-Model, we perform numerical simulations to investigate the emergence of collective slip dynamics for a given set of physical parameters.
 The simulations are carried out under prescribed frictional and elastic conditions, specified by the friction coefficient $\mu$, the spring constants $k_{p}$, and the number of blocks $N$.
 At each time step, the instantaneous force acting on each block is computed using Eqns.
-(11) and (12), and the corresponding Lorenz curve is constructed by plotting the force magnitude along the x-axis, and the cumulative fraction of blocks along the y-axis.
+(11) and (12), and the corresponding Lorenz curve is constructed by plotting the force magnitude along the y-axis, and the cumulative fraction of blocks along the x-axis.
 The degree of heterogeneity in the force distribution is quantified using inequality indices, which serve as compact descriptors of the non-uniformity in the stress landscape and provide a time-resolved measure of the buildup and release of inequality within the block ensemble.
@@
 For each value of the exchange range parameter ($R$) used in the B-Model, we calculate all three indices and present the results in Table 1.
-Additionally, motivated from a recent work [asim_2025], we examine the ratios p/(2k1)p/(2k-1) and $p/g$ to understand how these derived quantities evolve with the $g$ and $k$, offering further insights into the internal consistency and comparative behavior of these inequality measures.
+Additionally, motivated from a recent work [19], we examine the ratios p/(2k1)p/(2k-1) and $p/g$ to understand how these derived quantities evolve with the $g$ and $k$, offering further insights into the internal consistency and comparative behavior of these inequality measures.
 $R$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
@@
 Table 1: B-Model with Exchange Range: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for B-Model with different range $R$ computed from the Eqn (5).
-Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for B-model with different range ($R$).
+Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [19]) and $p/g$ values for B-model with different range ($R$).
 Figure 2 provides a visual summary of how the inequality measures evolve as we vary the range parameter in the B-Model.
@@
 In the upper-right panel, we illustrate the joint behaviors of the $g$-$k$ and $g$-$p$ pairs.
-As $R$ increases, $g$ and $k$ indices initially change in an almost perfectly linear manner with a slope $(3/8)$ as discussed in [bijin].
+As $R$ increases, $g$ and $k$ indices initially change in an almost perfectly linear manner with a slope $(3/8)$ as discussed in [22].
 During this early phase, the values of $g$ and $k$ coincide at a point approximately $0.85$.
@@
 Throughout the range, the ratio remains slightly above unity, reaching its maximum about a $3\%$ elevation near the point where $g$ and $k$ coincide (in this case, $g=k\sim 0.85$).
-This location has previously been identified as a precursor to criticality in several self-organized critical models [sand, manna, soumya].
+This location has previously been identified as a precursor to criticality in several self-organized critical models [2, 25, 15].
 The mild, non-monotonic rise toward this peak followed by a gradual decrease reflects a subtle shift in the balance between the $p$ and $k$ indices as the underlying wealth-exchange dynamics approach this near-critical region.
@@
 (3.1.2).
-Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for B-model with different saving propensities $\lambda$.
+Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [19]) and $p/g$ values for B-model with different saving propensities $\lambda$.
 In the upper-left panel of Figure 3, we present the relationship between $g$, $p$ and $k$ as the saving propensity $\lambda$ varies.
@@
 (7).
-Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for C-model with different saving propensities $\lambda$.
+Also showing the computed ratios p/(2k1)p/(2k-1) (motivated by [19]) and $p/g$ values for C-model with different saving propensities $\lambda$.
 The upper-left panel of Figure 4 illustrates how $g$, $p$ and $k$ co-varies with the saving propensity $\lambda$ in the C-Model or Yard-Sale Model framework.
@@
 As $\alpha$ approaches its lower bound of $1^{+}$, the distribution enters the regime of maximal permissible inequality while still maintaining a finite mean.
-Correspondingly, the indices converge toward their characteristic limiting values of approximately $0.85$.
+Correspondingly, the indices converge toward their characteristic limiting values.
 A decrease in $\alpha$ enhances the heaviness of the power-law tail, implying an increasingly concentrated allocation of income among a diminishing fraction of individuals.
@@
 Table 4: P-Model: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for P-model with different exponent $\alpha$, calculated using Eqn.
-(9), along with the computed ratios p/(2k1)p/(2k-1) (motivated by [asim_2025]) and $p/g$ values for P-model with different exponent $\alpha$.
+(9), along with the computed ratios p/(2k1)p/(2k-1) (motivated by [19]) and $p/g$ values for P-model with different exponent $\alpha$.
 The upper-left panel of Figure 5 shows how the inequality measures $g$, $p$, and $k$ evolve with the exponent $\alpha$.
@@
 Here too, we calculate $g$, $k$ and $p$, along with the ratios p/(2k1)p/(2k-1) and $p/g$, to understand how inequality evolves with each fractal generation.
-The table 4 shows values at 5-generation intervals from generation 5 onward, with generation 4 included separately.
-.
+The table 4 shows values at 5-generation intervals from generation 5 onward, with generations 4 and 199 included separately.
 $n$ $g$ $p$ $k$ p/(2k1)p/(2k-1) $p/g$
@@
 Table 5: CS-Model: Numerically estimated values of Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices for CS-model with different generations ($n$), where the overlap magnitude $Y_{n}(t)$ is calculated using Eqn.
-(10) along with the computed ratios p/(2k1)p/(2k-1) and $p/g$ (motivated by [asim_2025]).
+(10) along with the computed ratios p/(2k1)p/(2k-1) and $p/g$ (motivated by [19]).
 The upper-left panel of Figure 6 shows how the inequality measures $g$, $p$, and $k$ evolve with the generation number in the CS model.
@@
 Refer to captionRefer to caption
-Figure 7: BK-model: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with number of blocks (NN) = $100$, $\alpha=1000$, $k_{c}/k_{p}=1$ and $\sigma=0.01$ for BK-Model.
+Figure 7: BK-model: Results of numerical analysis of the variations of inequality indices Gini ($g$), Pietra ($p$) and Kolkata ($k$) indices with number of blocks (NN) = $100$, $\delta=1000$, $k_{c}/k_{p}=1$ and $\sigma=0.01$ for BK-Model.
 (a) Elastic energy released with respect to time ($t$); (b) $g$ vs $k$-index curve showing initial straight line fitting with slope $3/8$ and also $g$ vs $p$-index curve; (c) pp/(2k1)(2k-1) vs.
@@
 Surprisingly, in this regime of the model, $k$ and $g$ closely follow a linear relationship with slope $3/8$, similar to the previous examples.
-In the bottom left panel we show p/(2k1)p/(2k-1), where the values drop below 1, ranging from $\approx 0.98$ to $\approx 1.08$İn the bottom right panel, we plot $p/g$ after the initial transient period (first 100 steps) has passed and a steady state is reached, since the system starts.
+In the bottom left panel we show p/(2k1)p/(2k-1), where the values drop below 1, ranging from $\approx 0.98$ to $\approx 1.08$ .
+In the bottom right panel, we plot $p/g$ after the initial transient period (first 100 steps) has passed and a steady state is reached, since the system starts.
 In the Burridge-Knopoff model, the initial evolution is dominated by stress-loading transients arising from the imposed initial conditions.
@@
 In the bottom right panel, the ratio $p/g$ ranges from $\approx 0.69$ and to $\approx 0.82$
-6 Summery and Concluding remarks
-In this study, we have performed a comparative numerical investigation of inequality measures (namely of Gini [gini], Pietra [pietra] and Kolkata [k-index1, k-index3] indices) across a broad spectrum of model systems: they range from two kinetic wealth-exchange models (namely the Banerjee Model [B_Model, B_Model1] and Chakraborti model [C_Model], more popularly known as Yard-Sale model [YS, YS1]), to synthetic Pareto model [Pareto, Pareto1] and two earthquake-generating models (namely the Chakrabatrti-Stinchcombe model [CS_Model, CS1, pbhattacharya] and the Burridge-Knopoff model [langer_1994, BK_Model]).
+6 Summary and concluding remarks
+In this study, we have performed a comparative numerical investigation of inequality measures (namely of Gini [21], Pietra [29] and Kolkata [14, 3] indices) across a broad spectrum of model systems: they range from two kinetic wealth-exchange models (namely the Banerjee Model [5, 18] and
+Chakraborti model [12], more popularly known as Yard-Sale model [8, 7]), to synthetic Pareto model [28, 26] and two earthquake-generating models (namely the Chakrabatrti Stinchcombe model [10, 23, 6] and the Burridge Knopoff model [9, 27]).
 Using the Lorenz curves (see Fig.
@@
 2-7.
-It may be noted that while switching from the econophysical wealth exchange models (B and C models) which have got dominantly exponential (to Gamma-like) distributions and the two earthquake simulating models (CS and BK models) which have got (Guttenberg-Richter like) dominantly power law distributions, we considered and compared also the results for the perfect power-law Pareto distribution as well.
-We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models, the two-fractal model and the Pareto distributions) remain a little above unity, which was predicted to be exactly equal to unity [asim_2025].
-The value of this quantity for the two socio-economic (B and C models) and the earthquake simulating CS model is seen to deviate a little above unity, bounded by a maximum of $4\%$ near $g=k=0.86\pm 0.02$ which was identified earlier [manna, soumya] to be the precursor point of criticality in several self-organized critical models.
+It may be noted that while switching from the econophysical wealth exchange models (B and C models) which have got dominantly exponential (to Gamma-like) distributions and the two earthquake simulating models (CS and BK models) which have got (Gutenberg-Richter like) dominantly power law distributions, we considered and compared also the results for the perfect power-law Pareto distribution as well.
+We find, the values of p/(2k1)p/(2k-1) (across the wealth exchange models, the two-fractal model and the Pareto distributions) remain a little above unity, which was predicted to be exactly equal to unity [19].
+The value of this quantity for the two socio-economic (B and C models) and the earthquake simulating CS model is seen to deviate a little above unity, bounded by a maximum of $4\%$ near $g=k=0.86\pm 0.02$ which was identified earlier [25, 15] to be the precursor point of criticality in several self-organized critical models.
 For the Pareto distribution, the ratio p/(2k1)p/(2k-1) deviates slightly above unity, reaching a maximum of about $7\%$ (see Fig.
 5(c)).
-In the BK-Model, however, the value of p/(2k1)p/(2k-1) even drops below unity in certain regimes (see Fig.
+In the BK-Model, however, the value of p/(2k1)p/(2k-1) reaches a maximum of about $8\%$ above unity (see Fig.
 7(c)).
-The predicted values (assuming a simple polynomial form for the Lorenz function [asim_2025]) of $p/g=3/4$ seems to deviate quite a bit (in the range $0.69$ to about $1.0$; see Tables 1-5), while the existence of the relation $k=1/2+(3/8)g$ [asim_2025] seems to be valid for all the models considered here in the region of lower $g$ values (see Figs.
+The predicted values (assuming a simple polynomial form for the Lorenz function [19]) of $p/g=3/4$ seems to deviate quite a bit (in the range $0.69$ to about $1.0$; see Tables 1-5), while the existence of the relation $k=1/2+(3/8)g$ [19] seems to be valid for all the models considered here in the region of lower $g$ values (see Figs.
 2(d) - 7(d)).